OpenStudy (anonymous):

What is the simplified form of sin(x + p)? a. cos x b. sin x c. –sin x d. –cos x

ganeshie8 (ganeshie8):

you may use below formula : \(\large \sin (A + B) = \sin(A) \cos(B) + \cos(A) \sin (B)\)

ganeshie8 (ganeshie8):

seen that formula before ? :)

terenzreignz (terenzreignz):

Look what the wind dragged in... looks like trigonometry woes :3 Still... you might want to make your presence felt there ^^

ganeshie8 (ganeshie8):

looks you're scared and ran away after seeing that evil formula :o come back there is an easy way to do it :P

terenzreignz (terenzreignz):

Onde está você? <samba> :3

OpenStudy (anonymous):

Hi guysh... >.< I've never seen that formula before... lol ;-; Um... Does it look like this? sin(x) cos(p) + cos(x) sin(p) Why is Spanish being spoken lml >.<

terenzreignz (terenzreignz):

Não é espanhol, é português. Still though, does p mean \(\pi \) ?

ganeshie8 (ganeshie8):

okay im convinced... u didnt realize yet that u did ate some "pi" ? it should be sin(x + pi) right ?

terenzreignz (terenzreignz):

And lol, you plugged it in correctly, that's good ^^

OpenStudy (anonymous):

o.o I'm going to pretend I didn't see that xD Um... it is a possibility but it looks exactly like this What is the simplified form of sin(x + p)? a. cos x b. sin x c. –sin x d. –cos x It could be a mistake though... I've encountered so many mistakes throughout my course it's ridiculous. But it just says 'p' and Ooooh yay c:

OpenStudy (anonymous):

I don't like pi...e lol

terenzreignz (terenzreignz):

The only way the choices would make sense is if p were \(\large \pi\) Wouldn't you agree, ganeshie? :3

ganeshie8 (ganeshie8):

yess +1

terenzreignz (terenzreignz):

Yeah... well, the only thing to do now, Angel (can I call you Angel?) is recall what exactly these values are: \[\Large \sin(\pi) = \color{red}?\\\Large \cos(\pi ) = \color{red}?\]

terenzreignz (terenzreignz):

Stuck? Perhaps recalling the \(\large \pi\) corresponds to \(\large 180^o\) would help? ^^

terenzreignz (terenzreignz):

Uhh.... now's not a good time to space out @AngelCriner :3

OpenStudy (anonymous):

Sorry...... I was on Tumblr and the Cinnamon Rolls looked too amazing... I wanna cry ;-; Yesh of course you may lol But.. I'm a bit lost now >.<

OpenStudy (anonymous):

Do I just plug that into the calculator?

terenzreignz (terenzreignz):

Who needs to plug in? lol Well, if you haven't memorised it by now, you'd do well to start. \[\Large \sin(\pi) = 0\]\[\Large \cos(\pi ) = -1\]

OpenStudy (anonymous):

Does that turn into sin(x) (-1) + cos(x) (0)

terenzreignz (terenzreignz):

That it does :P So finally?

terenzreignz (terenzreignz):

Angel... <snap> <snap> You're almost there... don't die on me now ;)

OpenStudy (anonymous):

lol ok so p= pi :3

OpenStudy (anonymous):

Ooooh yay SORRY omg >.< >-\ Um.. I plugged sin(x) into mathway since I don't have a calculator and it just came out to be... sin(x) I don't think that's right e.e unless I wasn't supposed to do that...

terenzreignz (terenzreignz):

Don't worry about sin(x) Nobody knows what it is, just leave it be. So...back to simplifying? sin(x)(-1) + cos(x)(0)

OpenStudy (anonymous):

sin -1x + cos x

terenzreignz (terenzreignz):

lol what? Nothing goes inside the sin() -1 is multiplied to the whole sin(x) :P

terenzreignz (terenzreignz):

\[\Large \sin(x+\pi) = (\sin x) (\cos \pi)+ (\cos x)(\sin \pi)\]

OpenStudy (anonymous):

deleted lol it's ok ummmm I was told this sin(x) (-1) + cos(x) (0) -sin(x) + 0 -sin(x) but where does the -1 go? I'm not seeing it...

OpenStudy (anonymous):

OH WAIT

terenzreignz (terenzreignz):

Waiting for you to arrive at your realisation...

OpenStudy (anonymous):

since sin was positive and it's now negative... is that where it went o.o

OpenStudy (anonymous):

sin(x)

terenzreignz (terenzreignz):

Yup. Precisely. Very perceptive ^^

OpenStudy (anonymous):

Eh... kinda slow actually lol... thank you >.<

terenzreignz (terenzreignz):

So...any more trig you got?

OpenStudy (anonymous):

If you don't mind helping I can tag you in a couple of posts lol

terenzreignz (terenzreignz):

I'm bored... and we worked on permutations and combinations for about an hour and a half last night... do I look like I mind helping? XD

OpenStudy (anonymous):

XD I suppose... Point taken... Glad I can assist you in your boredom lol thank youuuu